OpenStudy (anonymous):

simplify 3×√(128a^13b^6

OpenStudy (coldplay):

Do you want us to simplify this?

OpenStudy (anonymous):

yes

OpenStudy (gorv):

\[3*\sqrt{128*a^{13}*b^6}\]

OpenStudy (gorv):

128=2^7

OpenStudy (gorv):

\[3*\sqrt{2^7*a^{13}*b^6}=3\sqrt{2*2^6*a*a^{12}*b^6}\]

OpenStudy (gorv):

can u do it now?

OpenStudy (coldplay):

Nope. Its too hard. Don't even judge me on it okay?